# Layer-wise Update Aggregation with Recycling for Communication-Efficient Federated Learning

### v1

@@ Abstract
Expensive communication cost is a common performance bottleneck in Federated Learning (FL), which makes it less appealing in real-world applications. Many communication-efficient FL methods focus on discarding a part of model updates mostly based on gradient magnitude. In this study, we find that recycling previous updates, rather than simply dropping them, more effectively reduces the communication cost while maintaining FL performance. We propose FedLUAR, a Layer-wise Update Aggregation with Recycling scheme for communication-efficient FL. We first define a useful metric that quantifies the extent to which the aggregated gradients influences the model parameter values in each layer. FedLUAR selects a few layers based on the metric and recycles their previous updates on the server side. Our extensive empirical study demonstrates that the update recycling scheme significantly reduces the communication cost while maintaining model accuracy. For example, our method achieves nearly the same AG News accuracy as FedAvg, while reducing the communication cost to just $17\%$.
@@ 1 Introduction
In this paper, we propose FedLUAR, a Layer-wise Update Aggregation with Recycling for communication-efficient and accurate FL. Instead of dropping the updates for a part of model parameters, we consider ‘ reusing’ the old updates multiple times in a layer-wise manner. Our study first defines a useful metric that quantifies the extent to which the aggregated gradients influences the model parameter values in each layer. Based on the metric, a small number of layers are selected to recycle their previous updates on the server side. Clients can omit these updates when sending their locally accumulated updates to the server. This layer-wise update recycling method allows only a subset of less important layers to lose their update quality while maintaining high-quality of updates for all the other layers.
Our study provides critical insights into achieving a practical trade-off between communication cost reduction and the level of noise introduced by any type of communication-efficient FL methods. First, by introducing noise into layers where the update magnitude is small relative to the model parameter magnitude, the adverse impact of the noise can be minimized, thus preserving FL performance. Second, our study empirically demonstrates that the update recycling approach achieves faster loss convergence compared to simply dropping updates for the same layers. We also theoretically analyze that FedLUAR converges to a neighborhood of the minimum when updates are recycled in a sufficiently small number of layers.
We evaluate the performance of FedLUAR using representative benchmark datasets: CIFAR-10 Krizhevsky (2009) , CIFAR-100, FEMNIST Caldas et al. (2018) , and AG News Zhang et al. (2015) We first compare FedLUAR to several state-of-the-art communication-efficient FL methods: Look-back Gradient Multiplier Azam et al. (2021), FedPAQ Reisizadeh et al. (2020), FedPara Hyeon-Woo et al. (2021), PruneFL Jiang et al. (2022), and FedDropoutAvg Gunesli et al. (2021). Then, we also compare the performance of advanced FL optimizers with and without our recycling method applied. Finally, we provide extensive ablation study results that further validate the efficacy of our proposed method, including performance comparisons based on the number of layers with recycled updates. These experimental results and our analysis show that FedLUAR provides a novel and efficient approach to reducing the communication cost while maintaining the model accuracy in FL environments.
@@ 2 Related Work
Structured Model Compression – Several FL methods based on low-rank decomposition have been studied, which re-parameterize the model parameters to save either computational cost or communication cost Phan et al. (2020); Vogels et al. (2020); Hyeon-Woo et al. (2021). These methods change the model architecture in a structured manner using various tensor approximation techniques Konečnỳ et al. (2016). The re-parameterization methods commonly increase the number of network layers, leading to higher implementation complexity as well as increased computational cost. Moreover, they struggle to maintain model performance when the rank is significantly reduced.
@@
Sketched Model Compression – Quantization-based FL methods have been actively studied, which reduce the number of bits per parameter Wen et al. (2017); Reisizadeh et al. (2020); Haddadpour et al. (2021); Chen and Vikalo (2024). Model pruning is another research direction which drops a part of model parameters to save not only computational cost but also communication cost Jiang et al. (2022, 2023); Li et al. (2024). These methods are categorized under sketched approach. Although quantization methods reduce communication overhead, they uniformly degrade the data representation quality of all parameters, disregarding their varying contributions to the training process. The pruning methods potentially harm the model’s learning capability since they directly reduce the number of parameters.
@@
Other Communication-Efficient FL Methods – FedLAMA has been proposed recently, which adaptively adjusts the model aggregation frequency in a layer-wise manner Lee et al. (2023b). Knowledge Distillation-based FL has also been studied Wu et al. (2022). These methods address the issue of high communication costs in FL; however, they do not consider the possibility of ‘reusing’ previous gradients. In this work, we focus on the potential to recycle previously computed gradients to accelerate communication-efficient FL.
@@
Gradient-Weight Ratio in Deep Learning – A few of recent works focus on utilizing gradient-weight ratio. Some researchers adjust learning rate based on the ratio to improve the model performance You et al. (2017); Mehmeti-Göpel and Wand (2024). These previous works theoretically demonstrate that the gradient-weight ratio deliver useful insights that can be utilized to adjust the inherent noise scale of stochastic gradients. In this study, we propose and employ a similar metric in FL environments: the ratio of accumulated updates to the initial model parameters at each communication round.
@@
Gradient Dynamics – Depending on the geometry of the parameter space, the gradient may remain consistent over several training iterations Azam et al. (2021). It has also been shown that the loss landscape becomes smoother as the batch size increases, and thus the stochastic gradients can remain similar for more iterations Keskar et al. (2017); Lin et al. (2020); Lee et al. (2023a). We explore the possibility of ‘recycling’ such stable gradients multiple times. By recycling previous updates, clients can avoid update aggregations in some network layers. In the following section, we will discuss how to safely recycle updates in a layer-wise manner.
@@ 3 Method
In this section, we first introduce a layer prioritization metric that can be efficiently calculated during training. Then, we present a communication-efficient FL method that recycle updates for layers with low priority. Finally, we provide a theoretical guarantee of convergence for the proposed FL method.
Figure 1: The layer-wise gradient norm and weight norm comparison (left) and the ratio of the gradient norm to the weight norm (right). The top-side (a) shows FEMNIST (CNN) and the bottom-side, (b) shows CIFAR-10 (ResNet20). It is clearly shown that the layers with the smallest gradients do not always least significantly affect the model parameter values.
@@ 3.1 Motivation
Many existing communication-efficient FL methods focus on how to reduce the communication cost while maintaining the gradient magnitude as similar as possible. For instance, update sparsification methods select a subset of parameters with small gradients and omit their updates Alistarh et al. (2018); Bibikar et al. (2021). Layer-wise model aggregation methods selectively aggregate the local updates at a subset of layers with small gradient norms Azam et al. (2021); Lee et al. (2023b). These methods commonly assume that larger gradients indicate greater importance.
Figure 1 shows layer-wise intermediate data collected from FEMNIST (top) and CIFAR-10 (bottom) training. The left-side chart compares the gradient norm and the weight norm while the right-side chart shows the ratio of the gradient norm to the weight norm. The IDs of a few layers with small gradient magnitudes (left) and ratios (right) are shown at the bottom of the charts. The detailed hyper-parameter settings can be found in Appendix.
@@
4:$\textbf{r}_{t}=[\hat{\Delta}_{t-1,l}],\forall l\in\mathcal{R}_{t}$
5:$\hat{\Delta}_{t}=[\textbf{r}_{t},\textbf{u}_{t}]$
6:Update the recycling scores: $\mathbf{s}_{t,l}$triangleright Eq. (1)
7:$\textbf{p}^{t}\leftarrow\text{Calculate }p_{l}^{t}$, $\forall l\in[L]$triangleright Eq. (2)
8:$\mathcal{R}_{t+1}\leftarrow$ Random_Choice($[L]$, $\delta$, $\textbf{p}^{t}$)
@@
3.2 Layer-wise Update Recycling
where $\Delta_{t,l}$ is the accumulated local updates averaged across all the clients at round $t$ for layer $l$, and $\mathbf{x}_{t,l}$ is the initial model parameters of layer $l$ at round $t$. Intuitively, this metric quantifies the relative gradient magnitude based on parameter magnitude. If $s_{t,l}$ is measured large, we expect the layer’s parameters to move fast on the parameter space making it sensitive to the update correctness. In contrast, if $s_{t,l}$ is small, the layer’s parameters will not be dramatically changed after each update. We assign low priority to layer $l$ if its $s_{t,l}$ is small, and high priority if it is large. In this way, all the $L$ layers can be prioritized based on how actively the parameters are changed after each round.
This metric can be efficiently measured at the server-side. The $\mathbf{x}_{t,l}$ is already stored in the server before every communication round. All FL methods aggregate the local updates after every round, and thus $\Delta_{t,l}$ is also already ready to be used in the server. Therefore, $s_{t,l}$ can be easily measured without any extra communications. This is a critical advantage considering the limited network bandwidth in typical FL environments.
Each layer has a weight factor $\frac{1}{s_{t,l}}$ so that it is less likely chosen if its priority is low. Dividing it by $\sum_{i=0}^{L-1}1/s_{t,l}$ ensures the sum of all weight factors equals $1$, allowing $p$ values to be directly used as weight factor of random sampling. Second, our method randomly samples $\delta$ layers using the probability distribution $\mathbf{p}$ shown in (2). We define those sampled layers at round $t$ as $\mathcal{R}_{t}$. Finally, the sampled $\delta$ layers are updated using the previous round’s updates instead of the latest updates. That is, the clients do not send the local updates for those $\delta$ layers to the server.
It is worth noting that the weighted random sampling-based layer selection prevents the updates for low-priority layers from being recycled excessively. When low-priority layers are not sampled, their updates will be normally aggregated at the server-side and thus their $s_{t,l}$ values can be updated. We will analyze the impact of this stochastic layer selection scheme on the overall performance of the update recycling method in Section 4.
@@ 3.2 Layer-wise Update Recycling
where ${\textbf{u}_{t}}$ is the updates for layers not included in $\mathcal{R}_{t}$, $\textbf{r}_{t}$ is the recycled updates for $\delta$ layers in $\mathcal{R}_{t}$, and $\hat{\Delta}_{t}$ is the global update composed of $\mathbf{u}_{t}$ and $\mathbf{r}_{t}$. Algorithm 1 shows Layer-wise Update Aggregation with Recycling (LUAR) method. Note that the number of layers whose update will be recycled, $\delta$, is a user-tunable hyper-parameter. We will further discuss how $\delta$ affects the model accuracy as well as the communication cost in Section 4.
Federated Learning Framework – Algorithm 2 shows FedLUAR, a FL framework built upon the proposed LUAR method shown in Algorithm 1. Before the active clients download the initial model parameters $\mathbf{x}_{t}$, the server lets them know the layers whose update will be recycled, $\mathcal{R}_{t}$ (line 5). Next, all individual clients run local training for $\tau$ iterations. Then, the local updates are aggregated using LUAR (line 11). Finally, the global model is updated using $\hat{\Delta}_{t}$. Figure 2 shows a schematic illustration of FedLUAR. If $\delta=0$, then $\hat{\Delta}_{t}$ becomes the same as $\Delta_{t}$, making it the vanilla FedAvg. In this paper, we use FedAvg as the base federated optimization algorithm for simplicity. However, extending it to more advanced FL optimizers is straightforward since LUAR is not dependent on any specific optimizer.
Potential Limitations – The proposed FL method consumes slightly increased memory space at the server-side as compared to the vanilla FedAvg. Specifically, as the server keeps $\delta$ selected layers’ previous updates, it requires a larger amount of memory space accordingly. In addition, the server should inform active clients about which layers’ updates should be omitted when transmitting the local updates to the server. This introduces only a negligible overhead, as the list of recycling layer IDs can be sent along with the initial model parameters before each communication round.
Figure 2: A schematic illustration of FedLUAR. Each client sends out updates for layers with large $s_{t,l}$ values only. For the other layers, the server ‘recycles’ the previous updates.
@@ 3.3 Theoretical Analysis
where $F_{i}(\mathbf{x})=\mathbb{E}_{\xi_{i}\sim D_{i}}[f(\mathbf{x},\xi_{i})]$ is the local loss function associated with the local data distribution $D_{i}$ of client $i$ and $m$ is the number of clients.
Our analysis is based on the following assumptions.
@@
We define $\hat{g}$ as a stochastic gradient vector that has zero in the selected $\delta$ layers where their updates will be recycled. Likewise, the corresponding full-batch gradient is defined as $\nabla\hat{F}(\mathbf{x})$. To analyze the impact of the update recycling in Algorithm 1, we also define the quantity of noise $n_{t}$ as follows.
@@
Herein, we present our theoretical analysis on the convergence rate of Algorithm 2 (See Appendix for proofs).
Lemma 3.1.
@@
where $m$ is the number of clients and $\kappa$ is the ratio of $\|\nabla\hat{F}(\mathbf{x}_{t})\|^{2}$ to $\|\nabla F(\mathbf{x}_{t})\|^{2}$ which is $\leq 1$.
Theorem 3.2.
@@
Remark 1. Lemma 3.1 shows that the update recycling method ensures the noise magnitude bounded regardless of how many times the updates are recycled and how many layers recycle their updates if $\kappa$ is sufficiently small. This result can serve as a foundation that allows users to safely recycle updates in a layer-wise manner, thereby reducing communication costs.
Remark 2. Algorithm 2 converges to a neighbor region of a minimum since the term of $(4+9\mathcal{L}^{2})\sigma_{L}^{2}$ in (7) is not affected by the learning rate $\eta$. However, all the other terms are zeroed out as $\eta$ diminishes. Thus, we can conclude that Algorithm 2 ensures practical convergence properties in real-world applications.
Intuitively, Theorem 3.2 shows that the data heterogeneity affects how closely the model converges to a minimum. To mitigate this effect, one could use a relatively large batch size during local training steps to help the model converge as closely as possible to the minimum.
Algorithm CIFAR-10 CIFAR-100 FEMNIST AG News

### v2

@@ Abstract
Expensive communication cost is a common performance bottleneck in Federated Learning (FL), which makes it less appealing in real-world applications. Many communication-efficient FL methods focus on discarding a part of model updates mostly based on gradient magnitude. In this study, we find that recycling previous updates, rather than simply dropping them, more effectively reduces the communication cost while maintaining FL performance. We propose FedLUAR, a Layer-wise Update Aggregation with Recycling scheme for communication-efficient FL. We first define a useful metric that quantifies the extent to which the aggregated gradients influence the model parameter values in each layer. FedLUAR selects a few layers based on the metric and recycles their previous updates on the server side. Our extensive empirical study demonstrates that the update recycling scheme significantly reduces the communication cost while maintaining model accuracy. For example, our method achieves nearly the same AG News accuracy as FedAvg, while reducing the communication cost to just $17\%$.
@@ 1 Introduction
In this paper, we propose FedLUAR, a Layer-wise Update Aggregation with Recycling method for communication-efficient and accurate FL. Instead of dropping the updates for a part of model parameters, we consider ‘reusing’ the old updates multiple times in a layer-wise manner. Our study first defines a useful metric that quantifies the extent to which the aggregated gradients influence the model parameter values in each layer. Based on the metric, a small number of layers are selected to recycle their previous updates on the server side. Clients can omit these updates when sending their locally accumulated updates to the server. This layer-wise update recycling method allows only a subset of less important layers to lose their update quality while maintaining high-quality updates for all the other layers. Our study shows that, by carefully choosing the update recycling layers, the model aggregation cost can be dramatically reduced while maintaining the model accuracy.
Our study provides critical insights into achieving a practical trade-off between communication cost reduction and the level of noise introduced by any types of communication-efficient FL methods. First, by introducing noise into layers where the update magnitude is small relative to the model parameter magnitude, the adverse impact of the noise can be minimized, thus preserving FL performance. Second, our study empirically demonstrates that the update recycling approach achieves faster loss convergence compared to simply dropping updates for the same layers. Our theoretical analysis also shows that FedLUAR converges to a neighborhood of a stationary point when updates are recycled in a sufficiently small number of layers. We designed our FL method to leverage these findings and it can be readily applied to various FL applications to improve scalability.
We evaluate the performance of FedLUAR  [^1] https://github.com/swblaster/FedLUAR using representative benchmark datasets: CIFAR-10 [19], CIFAR-100, FEMNIST [4], and AG News [42] We first compare FedLUAR to several state-of-the-art communication-efficient FL methods: Look-back Gradient Multiplier [2], FedPAQ [31], FedPara [12], PruneFL [13], FedDropoutAvg [8], and FedBAT [23]. We also compare the performance between using and not using the recycling method for advanced FL optimizers. Finally, we provide extensive ablation study results that further validate the efficacy of our proposed method, including performance comparisons based on the number of layers with recycled updates. These experimental results and our analysis show that FedLUAR provides a novel and efficient approach to reducing the communication cost while maintaining the model accuracy in FL environments.
@@ 2 Related Work
Structured Model Compression – Several low-rank decomposition-based FL methods have been proposed, which re-parameterize the model weights to reduce either computational or communication costs [28, 35, 12]. These methods modify the model architecture in a structured way using various tensor approximation techniques [18]. The re-parameterization methods often increase the number of network layers, resulting in higher implementation complexity and higher computational costs. Moreover, they struggle to maintain model performance when the rank is significantly reduced.
@@
Sketched Model Compression – Quantization-based FL methods have been actively studied to reduce the number of bits used per parameter [36, 31, 9, 5]. Model pruning methods, such as PruneFL [13], FedMP [14], FedPruning [22], GossipFL [33], and SpaFL [17], remove a portion of model parameters to reduce both computational and communication costs. These methods are categorized under a sketching approach. Although quantization methods reduce communication overhead, they uniformly degrade the data representation quality of all parameters, overlooking their varying contributions to the training process. The pruning methods potentially harm the model’s learning capability since they directly reduce the number of parameters.
@@
Other Communication-Efficient FL Methods – FedLAMA [21] adaptively adjusts model aggregation frequency in a layer-wise manner. Dynamic model aggregation method proposed in [15] aggregates local models in a decentralized manner. FedKD [37] reduces communication cost by employing knowledge distillation in place of model aggregation. These methods address the high communication cost issue in FL. However, they do not consider the possibility of ‘reusing’ previous gradients. In this work, we focus on recycling previously computed gradients to reduce the communication cost. Bandwidth-aware Compression Ratio Scheduling (BCRS) [34] adjusts the top-k compression ratio in a network bandwith-aware manner. YOGA [26] adopts a layer-wise aggregation strategy based on layer priority, which shares conceptual similarities with our proposed method. However, it assumes a peer-to-peer decentralized FL environment without a central server, which is not applicable to server-based FL scenarios.
@@
Gradient-Weight Ratio in Deep Learning – A few of the recent works focus on utilizing gradient-weight ratio. Some researchers adjust learning rate based on the ratio to improve the model performance [39, 27]. These previous works theoretically demonstrate that the gradient-weight ratio delivers useful insights that can be utilized to adjust the inherent noise scale of stochastic gradients. In this study, we propose and employ a similar metric in FL environments: the ratio of accumulated updates to the initial model parameters at each communication round.
@@
Gradient Dynamics – Depending on the geometry of the parameter space, the gradient may remain consistent over several training iterations [2]. It has also been shown that the loss landscape becomes smoother as the batch size increases, and thus the stochastic gradients can remain similar for more iterations [16, 25, 20]. We explore the possibility of ‘recycling’ such stable gradients multiple times. By recycling previous updates, clients can avoid update aggregation in some network layers. In the following section, we will discuss how to safely recycle updates in a layer-wise manner.
@@ 3 Method
In this section, we first introduce a layer prioritization metric that can be efficiently calculated during training. Then, we present a communication-efficient FL method that recycles updates for layers with low priority. Finally, we provide a theoretical guarantee of convergence for the proposed FL method.
@@ 3.1 Motivation
(a) FEMNIST (CNN)
(b) CIFAR-10 (ResNet20)
Figure 1: The layer-wise gradient norm and weight norm comparison (left) and the ratio of the gradient norm to the weight norm (right). It is clearly shown that the layers with the smallest gradients do not always least significantly affect the model parameter values.
Many existing communication-efficient FL methods focus on how to reduce the communication cost while keeping the gradient magnitude as close as possible to the original. For instance, update sparsification methods select a subset of parameters with small gradients and omit their updates [1, 3]. Layer-wise model aggregation methods selectively aggregate the local updates at a subset of layers with small gradient norms [2, 21]. These methods commonly assume that larger gradients indicate greater importance.
Figure 1 shows layer-wise intermediate data collected from FEMNIST and CIFAR-10 training. The left-side chart compares the gradient norm and the weight norm while the right-side chart shows the ratio of the gradient norm to the weight norm. The IDs of a few layers with small gradient magnitudes (left) and ratios (right) are shown at the bottom of the charts. The detailed hyper-parameter settings can be found in Appendix.
@@
3.2 Layer-wise Update Recycling
where $\Delta_{t,l}$ is the accumulated local updates averaged across all the clients at round $t$ for layer $l$, and $\mathbf{x}_{t,l}$ is the initial model parameters of layer $l$ at round $t$. Intuitively, this metric quantifies the relative gradient magnitude based on parameter magnitude. If $s_{t,l}$ is measured large, we expect the layer’s parameters to move fast in the parameter space making it sensitive to the update correctness. In contrast, if $s_{t,l}$ is small, the layer’s parameters will not be dramatically changed after each update. We assign low priority to layer $l$ if its $s_{t,l}$ is small, and high priority if it is large. In this way, all the $L$ layers can be prioritized based on how actively the parameters are changed after each round.
This metric can be efficiently measured on the server-side. The $\mathbf{x}_{t,l}$ is already stored on the server before every communication round. All FL methods aggregate the local updates after every round, and thus $\Delta_{t,l}$ is also already ready to be used on the server. Therefore, $s_{t,l}$ can be easily measured without any extra communications. This is a critical advantage considering the limited network bandwidth in typical FL environments.
Each layer has a weight factor $\frac{1}{s_{t,l}}$ so that it is less likely chosen if its priority is low. Dividing it by $\sum_{i=0}^{L-1}1/s_{t,l}$ ensures the sum of all weight factors equals $1$, allowing $p$ values to be directly used as a weight factor of random sampling. Second, our method randomly samples $\delta$ layers using the probability distribution p shown in (2). We define those sampled layers at round $t$ as $\mathcal{R}_{t}$. Finally, the sampled $\delta$ layers are updated using the previous round’s updates instead of the latest updates. That is, the clients do not send to the server the local updates for those $\delta$ layers.
It is worth noting that the weighted random sampling-based layer selection prevents the updates for low-priority layers from being recycled excessively. When low-priority layers are not sampled, their updates will be normally aggregated on the server-side and thus their $s_{t,l}$ values can be updated. We will analyze the impact of this stochastic layer selection scheme on the overall performance of the update recycling method in Section 4.
@@ 3.2 Layer-wise Update Recycling
[…]
6:Update the recycling scores: $\mathbf{s}_{t,l}$ $\triangleright$ Eq. (1)
[…]
@@
where $\textbf{u}_{t}$} is the updates for layers not included in $\mathcal{R}_{t}$, $\textbf{r}_{t}$ is the recycled updates for $\delta$ layers in $\mathcal{R}_{t}$, and $\hat{\Delta}_{t}$ is the global update composed of $\mathbf{u}_{t}$ and $\mathbf{r}_{t}$. Algorithm 1 shows the Layer-wise Update Aggregation with Recycling (LUAR) method. Note that the number of layers whose update will be recycled, $\delta$, is a user-tunable hyper-parameter. We will further discuss how $\delta$ affects the model accuracy as well as the communication cost in Appendix A.4.
Figure 2: Schematic illustration of FedLUAR.
Federated Learning Framework – Algorithm 2 shows FedLUAR, a FL framework built upon LUAR (Alg. 1). Before the active clients download the initial model parameters $\mathbf{x}_{t}$, the server informs them of the set of layers to be recycled, denoted by $\mathcal{R}_{t}$ (line 5). Subsequently, each client performs local training for $\tau$ iterations. The resulting local updates are then aggregated using LUAR (line 11). Finally, the global model is updated using $\hat{\Delta}_{t}$. Figure 2 shows a schematic illustration of FedLUAR. Each client transmits updates only for the layers with large $s_{t,l}$ values. For the remaining layers, the server recycles the previous updates. When $\delta=0$, $\hat{\Delta}_{t}$ becomes the same as $\Delta_{t}$, effectively reducing the method to vanilla FedAvg. In this paper, we use FedAvg as the base federated optimization algorithm for simplicity. However, extending it to more advanced FL optimizers is straightforward, as LUAR is agnostic to the choice of optimizer.
Potential LimitationsFedLUAR requires the server to notify active clients about which layer updates should be omitted when uploading their local updates. This introduces an additional communication cost compared to other FL methods. However, this extra overhead is likely negligible, as the list of recycled layer IDs ($\delta$ integers) can be transmitted along with the initial model parameters at the beginning of each communication round.
@@ 3.3 Theoretical Analysis
where $F_{i}(\mathbf{x})=\mathbb{E}_{\xi_{i}\sim D_{i}}[f(\mathbf{x},\xi_{i})]$ is the local loss function associated with the local data distribution $D_{i}$ of client $i$ and $m$ is the number of clients. Our analysis is based on the following assumptions.
@@
Noise Definition – We define $\hat{g}$ as a stochastic gradient vector that corresponds to the $\delta$ layers whose updates will be recycled. Likewise, the corresponding full-batch gradient is defined as $\nabla\hat{F}(\mathbf{x})$. We quantify the ratio of $\|\nabla\hat{F}(\mathbf{x}_{t})\|^{2}$ to $\|\nabla F(\mathbf{x}_{t})\|^{2}$, which is $\leq 1$, as $\kappa$. By the definition of $\nabla\hat{F}(\mathbf{x})$, $\kappa$ goes to zero if none of the layers recycle their updates. To analyze the impact of the update recycling in Algorithm 1, we also define the quantity of noise $n_{t}$ as follows.
@@
Herein, we analyze the convergence rate of Algorithm 2 (See Appendix for proofs).
Lemma 1.
@@
where $m$ is the number of clients.
Theorem 2.
@@
Remark 1. Lemma 1 shows that the update recycling method ensures the noise magnitude bounded regardless of how many times the updates are recycled and how many layers recycle their updates if $\kappa$ is sufficiently small. This result can serve as a foundation that allows users to safely recycle updates in a layer-wise manner, thereby reducing communication costs.
Remark 2. Algorithm 2 converges to a neighborhood of a stationary point, as the second and third terms on the right-hand side of (7) do not vanish as $T\rightarrow\infty$. Furthermore, the term of $(4+9\mathcal{L}^{2})\sigma_{L}^{2}$ is independent of the learning rate $\eta$ and remains non-zero even as $\eta\rightarrow 0$. Although it does not ensure convergence to an exact minimum, this rough guarantee is considered useful in real-world applications [41, 7].
In general, as the degree of non-IIDness increases, the global variance tends to grow due to greater discrepancies among local datasets. As shown in the final term on the right-hand side of (7), recycling updates in more layers increases the coefficient $\frac{1080T\mathcal{L}^{2}\eta^{2}\tau^{2}}{1-16\kappa}$, which in turn amplifies the final term. Consequently, the model is expected to converge more slowly. This suggests using smaller learning rate as the degree of non-IIDness increases to maintain the convergence rate.
3.4 Memory Usage Analysis
Dataset (Model) Algorithm $\delta$ Memory
Footprint (MB)
CIFAR-10 (ResNet20) FedAvg - 33.49
FedLUAR 10 15.23
CIFAR-100 (WRN28-10) FedAvg - 4,462.80
FedLUAR 14 2,604.88
FEMNIST (CNN) FedAvg - 806.11
FedLUAR 2 204.73
AG News (DistillBERT) FedAvg - 8,294.18
FedLUAR 30 1,825.42
Table 1: Comparison of memory usage observed during training.
In FedAvg, server should receive local models from all active clients. Consequently, the maximum memory footprint is $a\cdot d$, where $a$ is the number of active clients and $d$ is the model size. By contrast, FedLUAR receives local models from active clients except $\delta$ layers. Thus, the memory footprint is $a\cdot(d-k)$, where $k$ is the size of $\delta$ layers. Instead, the previous global update should be kept in the memory space for the $\delta$ layers, consuming $k$ space only. Therefore, FedLUAR’s memory footprint is $a\cdot(d-k)+k<a\cdot d$.
To support our analysis, we actually measured the memory footprint of FedAvg and FedLUAR during training. First, the total number of clients is 128 and only randomly selected 32 clients are activated at each communication round. We use MPI to run FL on 2 GPUs. Thus, each process locally train 16 models and then all the locally trained models are aggregated using MPI_Allreduce(). Table 1 shows the memory footprint of each process, observed during training under this setting. We can clearly see that FedLUAR uses less memory space than FedAvg. This advantage is directly related to the reduced communication cost, which will be discussed in Section 4.3.
4 Experiments
Experimental Settings – All experiments are conducted on a GPU cluster which has 2 NVIDIA A6000 GPUs per machine. We use TensorFlow 2.15.0 for training and MPI for model aggregations. All experiments were performed at least 3 times, and the average accuracies are reported.
Datasets – We evaluate the performance of our proposed method on representative benchmarks: CIFAR-10 (ResNet20 [10]), CIFAR-100 (Wide-ResNet28 [40]), FEMNIST (CNN), and AG News (DistillBert [32]). When tuning hyper-parameters, we conduct a grid search with a sufficiently small unit size (e.g., 0.1 for learning rate). To generate non-IID datasets, we use label-based Dirichlet distributions with $\alpha=0.1$, which indicates highly non-IID conditions.
Data Heterogeneity – For IID datasets, we simulate non-IID settings using Dirichlet distributions. The concentration coefficient $\alpha$ is set to 0.1 for CIFAR-10/100 and 0.5 for AG News.
4.1 Comparative Study
We first present an accuracy comparison among SOTA communication-efficient FL methods below.
  • LBGM (Low-rank Approximation) [2]
  • FedPAQ (Quantization) [31]
  • FedPara (Reparameterization) [12]
  • PruneFL (Pruning) [13]
  • FedDropoutAvg (Dropping) [8]
  • FedBAT (Binarization) [23]
Table 2 shows the performance comparison (See Appendix for the detailed settings). The total number of clients is 128 and randomly chosen 32 clients participate in every communication round. Note that the FL methods cannot have exactly the same communication cost due to differences in their mechanisms. To ensure fair comparisons, we find algorithm-specific settings that achieve accuracy reasonably close to the baseline (FedAvg) while minimizing communication costs, and then compare the validation accuracy across algorithms.
Overall, FedLUAR achieves accuracy comparable to the baseline while significantly reducing communication costs across all four benchmarks. Notably, for FEMNIST and AG News, it matches FedAvg’s accuracy with less than $20\%$ of the communication cost. Our method also outperforms all other SOTA methods. While FedPAQ and FedBAT reduce communication cost, they suffer from noticeable accuracy drops. Regardless of the dataset, FedLUAR consistently delivers the highest accuracy among communication-efficient FL methods. These results demonstrate that LUAR effectively finds less critical layers and recycles their updates, minimizing the communication cost without sacrificing performance.
Method CIFAR-10 CIFAR-100 FEMNIST AG News